The dimensions of a rectangle garden are 3 feet by 10 feet. When both demensions are increased by the same amount, the area is doubled. What are the demensions of the new garden?
Answer by ankor@dixie-net.com(22740)

The dimensions of a rectangle garden are 3 feet by 10 feet.
When both dimensions are increased by the same amount, the area is doubled.
What are the dimensions of the new garden?
:
Find the original area: 3 * 10 = 30, then obviously 60 sq/ft is the doubled area
:
Let x = increase of length and width to double the area
(x+3)(x+10) = 60
FOIL
x^2 + 10x + 3x + 30 - 60 = 0
A quadratic equation
x^2 + 13x - 30 = 0
Factors easily to
(x+15)(x-2) = 0
Positive solution is all we want here
x = 2
New dimensions: 5 by 12 (which = 60 sq/ft)